Alexander paid $148,000 to acquire 100% of Willis Corporation in a statutory merger. Alexander also agreed to pay the shareholders of Willis $0.80 in cash for every dollar in income from continuing operations of the combined entity over $75,000 in the first year following acquisition. Alexander projects that there is a 10% (45%, 25%, 20%) probability that the income from continuing operations for the year is $65,000 ($75,000, $85,000, $95,000 respectively). Alexander uses a discount rate of 8%.

Information for Willis Corporation immediately before the merger was as follows:

Book ValueFair Value

Current Assets$40,000$50,000

Plant Assets$120,000$70,000

Liabilities$50,000$45,000

Previously unreported items identified as belonging to Willis:

Fair Value

Contracts under negotiation with potential customers$15,000

Customer contracts for consulting projects$7,000

In-process research and development$8,000

Skilled workforce$23,000

Recent favorable press reports on Willis$2,000

Proprietary databases$12,000

Determine the goodwill to be reported in this acquisition.
Answer #1

Value of goodwill = Consideration paid + Fair value of non controlling interest + Fair value of equity previous interests - Fair value of net assets recognized.

Consideration paid = 148,000$

Cash portion = 55000$ (W N - 1)

Fair value of non controlling interest = Nil (as 100% equity is purchased and there is no minority interest)

Fair value of equity previous interests = Nil (as no other equity interest is mentioned)

Fair value of net assets = 125,000 (W N - 2)

Applying above formula we get, Goodwill = 148000 + 55000 - 125000

= 78,000$

W N - 1

(Projected Earnings after year of operations (65000*10% + 75000*45% + 85000*25% + 95000*20%) = 80,500$

Projected earnings over 75000$ = 5500$

Discounting rate = 8%

Hence, projected earnings = 5500/0.08 = 68750 $

Cash portion = 80% of 68750 = 55000$)

W N - 2

Current Assets are resources owned by a company and which have future economic value that can be measured and can be expressed in dollars. Hence, Contracts under negotiation with potential customers/recent favorable reports cannot be considered assets as we dont have reasonable assurance that these have future economic value.  

Add Current Assets - 50,000$

Add Plant Assets - 70,000$

Add Customer contracts for consulting projects$7,000

Add In-process research and development $8,000

Add Skilled workforce $23,000

Add Proprietary databases$12,000

Less Liabilities - 45,000$

= 125000 $